Consider the boilers made up of different layers of conducting materials. How the heat conduction takes place in such an arrangement when they are connected in series and parallel?
1
Expert's answer
2020-12-14T12:17:04-0500

Calculate temperature difference on the opposite sides of conducting materials connected with thermal resistance r and R in series and parallel assuming that the power Q is equal for both cases:


"\Delta T_s=Q(r+R),\\\space\\
\Delta T_p=Q\bigg(\frac{rR}{r+R}\bigg)."

As we see, the temperature difference will be greater for series connection and smaller for parallel connection, which determines application of these such arrangements.